1. Cheraw Auto Repair believes that it will need new equipment in 10 years. The equipment will cost 11) $26,000. What lump sum should be invested today at 8% compounded semiannually, to yield $26,000?

2. 4x2 x 2 Find the values of x where the tangent line is horizontal for the graph of f(x)- A)x=-2 C)x=-2, x=0, x=-4 B)x=0,x=-2
Answer #1

Semi annual interest=(8/2)=4%=0.04

Future value (FV)=Present Value (PV)*((1+i)^N)

i=Interest per period=0.04

N=Number of Periods=10*2=20

Future Value (FV)=$26000

Lump sum to be invested today=PV=FV/((1+i)^N)=26000/(1.04^20)= $11,866.06

Lumpsum to be invested today

$11,866.06